# Local Advice and Local Decompression

### v1

@@
Abstract.
Algorithms with advice have received ample attention in the distributed and online settings, and they have recently proven useful also in dynamic settings. In this work we study local computation with advice: the goal is to solve a graph problem $\Pi$ with a distributed algorithm in $f(\Delta)$ communication rounds, for some function $f$ that only depends on the maximum degree $\Delta$ of the graph, and the key question is how many bits of advice per node are needed. Our main results are:
  1. (1)
    Any locally checkable labeling problem (LCL) can be solved in graphs with sub-exponential growth with only $1$ bit of advice per node. Moreover, we can make the set of nodes that carry advice bits arbitrarily sparse.
  2. (2)
    The assumption of sub-exponential growth is necessary: assuming the Exponential-Time Hypothesis (ETH), there are LCLs that cannot be solved in general with any constant number of bits per node.
  3. (3)
    In any graph we can find an almost-balanced orientation (indegrees and outdegrees differ by at most one) with $1$ bit of advice per node, and again we can make the advice arbitrarily sparse.
  4. (4)
    As a corollary, we can also compress an arbitrary subset of edges so that a node of degree $d$ stores only $d/2+2$ bits, and we can decompress it locally, in $f(\Delta)$ rounds.
  5. (5)
@@ Abstract.
  6. (6)
    In any $3$-colorable graph, we can find a $3$-coloring with $1$ bit of advice per node. Here, it remains open whether we can make the advice arbitrarily sparse.
Our work shows that for many problems the key threshold is not whether we can achieve, say, $1$ bit of advice per node, but whether we can make the advice arbitrarily sparse. To formalize this idea, we develop a general framework of composable schemas that enables us to build algorithms for local computation with advice in a modular fashion: once we have (1) a schema for solving $\Pi_{1}$ and (2) a schema for solving $\Pi_{2}$ assuming an oracle for $\Pi_{1}$, we can also compose them and obtain (3) a schema that solves $\Pi_{2}$ without the oracle. It turns out that many natural problems admit composable schemas, all of them can be solved with only $1$ bit of advice, and we can make the advice arbitrarily sparse.
This work was supported by the Independent Research Fund Denmark grant 2020-2023 (9131-00044B) “Dynamic Network Analysis” and the VILLUM Foundation grant (VIL37507) “Efficient Recomputations for Changeful Problems”
1. Introduction
@@
While computation with advice has been explored in a wide range of distributed settings (Fraigniaud et al., 2007, 2010; Dobrev et al., 2012; Glacet et al., 2017; Dereniowski and Pelc, 2012; Gorain and Pelc, 2019; Miller and Pelc, 2015a, 2016; Fusco and Pelc, 2011; Ilcinkas et al., 2010; Nisse and Soguet, 2009; Fraigniaud et al., 2008; Komm et al., 2015; Miller and Pelc, 2015b; Fusco et al., 2016), there is hardly any prior work on solving classic local graph problems. A rare example of prior work is (Fraigniaud et al., 2009) from 2007, which studied the question of how much advice is necessary to break Linial’s (Linial, 1992) lower bound for coloring cycles.
We initiate a systematic study of exactly how much advice is needed in the context of a wide range of graph problems. As we will see in this work, the exploration of the advice complexity of graph problems opens up connections with many other topics—it is linked with distributed proofs (Korman and Kutten, 2006, 2007; Korman et al., 2010a, b; Göös and Suomela, 2016; Feuilloley and Fraigniaud, 2017), distributed decompression, the notion of order-invariant algorithms (Naor and Stockmeyer, 1995), and also with the exponential-time hypothesis (Impagliazzo and Paturi, 1999) in computational complexity theory.
1.1. Local computation with advice
@@
Our work shows that for many problems the key threshold is not whether we can achieve, say, $1$ bit of advice per node, but whether we can make the advice arbitrarily sparse, that is, make the ratio between $1$s and 00s assigned to the nodes of the graph to be an arbitrarily small constant. As such, a decent part of this paper addresses the following questions.
Which problems can be solved with $1$ bit of advice distributed to an arbitrary sparse set of vertices? Are there any problems that cannot be solved with $1$ bit of advice?
@@
The notion of the sparsity of the advice is discussed later in the paper. In this context, our paper shows that some problems can be solved with arbitrarily sparse advice. On the other hand, we also show that assuming the Exponential-Time Hypothesis, for any constant $c$, there exist problems that cannot be solved with $c$ bits of advice. Finally, there are some problems like 3-coloring that can be solved with 1 bit of advice, but where it is not clear whether it can be solved with arbitrarily sparse advice. We discuss all of those points in more detail in the remaining part of this section.
1.2. Contribution 1: LCLs in bounded-growth graphs
@@
LCLs, first introduced by Naor and Stockmeyer (1995) in the 1990s, are one of the most extensively studied families of problems in the theory of distributed graph algorithms. These are graph problems that can be specified by giving a finite set of valid local neighborhoods. Many key problems such as vertex coloring, edge coloring, maximal independent set, maximal matching, sinkless orientation, and many other splitting and orientation problems are examples of LCLs, at least when restricted to bounded-degree graphs. Thanks to the extensive research effort since 2016, we now understand very well the landscape of all LCL problems and their computational complexities across different models of distributed computing (Balliu et al., 2018, 2021a; Chang and Pettie, 2019; Ghaffari et al., 2018; Balliu et al., 2020; Fischer and Ghaffari, 2017; Rozhon and Ghaffari, 2020; Brandt et al., 2016; Chang et al., 2019; Ghaffari and Su, 2017; Balliu et al., 2021b).
@@
In Section 4, we design a schema that allows us to solve any LCL problem with just one bit of advice in graphs with a sub-exponential growth (the number of nodes in a radius-$r$ neighborhood is sub-exponential in $r$):
@@
Any LCL problem can be solved with $1$ bit of advice per node in graphs with sub-exponential growth.
Furthermore, we show that the encoding can be made arbitrarily sparse.
@@
One prominent application of this result is its connection with distributed proofs (Korman and Kutten, 2006, 2007; Korman et al., 2010a, b; Feuilloley and Fraigniaud, 2017), and in particular with locally checkable proofs (Göös and Suomela, 2016). Consider any LCL $\Pi$. Assume that our task is to prepare a distributed proof that shows that in a graph $G$ there exists a feasible solution of $\Pi$. Now if $G$ has sub-exponential growth, we can use our result from Section 4 to prepare a $1$-bit advice that enables the algorithm to find a solution of $\Pi$. Our advice is the proof: to verify it, we simply try to recover a solution with the help of the advice, and then check that the output is feasible in all local neighborhoods (recall that $\Pi$ is locally checkable). We obtain the following corollary:
@@
Note that this is not a proof labeling scheme as defined in (Korman and Kutten, 2006, 2007; Korman et al., 2010a, b), as the verifier inspects a hop-neighborhood whose radius, albeit constant, may be more than $1$.
1.3. Contribution 2: LCLs in general graphs
At this point a natural question is whether the assumption about bounded growth is necessary. Could we solve all LCL problems in all graphs with $1$ bit of advice? In Section 8 we show that the answer is likely to be no:
@@ 1.3. Contribution 2: LCLs in general graphs
Hence, we need to show that $\mathcal{A}$ can be made cheap to simulate. The key ingredient is the following technical result, which we prove using a Ramsey-type argument that is inspired by the proof of Naor and Stockmeyer (1995):
@@
The key point here is that (for bounded-degree graphs) $\mathcal{A}^{\prime}$ can be represented as a finite lookup table; hence the simulation of $\mathcal{A}^{\prime}$ is cheap, and we can finally make a formal connection to the Exponential-Time Hypothesis.
1.4. Contribution 3: balanced orientations
In Section 5 we move on to a specific graph problem: we study the task of finding balanced and almost-balanced orientations. The goal is to orient the edges so that for each node indegree and outdegree differ by at most $1$. This is a hard problem to solve in a distributed setting, while slightly more relaxed versions of the problem admit efficient (but not constant-time) algorithms (Ghaffari et al., 2020).
@@ 1.4. Contribution 3: balanced orientations
1.5. Contribution 4: distributed decompression
Equipped with the advice schema for solving almost-balanced orientations, we can now make a formal connection to distributed decompression. Here the task is to encode some graph labeling so that it can be decompressed locally (in $T(\Delta)$ rounds).
@@ 1.5. Contribution 4: distributed decompression
1.6. Contribution 5: vertex $\Delta$-coloring
In Section 6 we study the problem of finding a $\Delta$-coloring in graphs of maximum degree $\Delta$. Our main result there is:
@@ 1.6. Contribution 5: vertex $\Delta$-coloring
Our schema for encoding $\Delta$-colorings consists of three steps. First, we compute a vertex coloring with $O(\Delta^{2})$ colors, with the help of advice. Then we reduce the number of colors down to $\Delta+1$, using the algorithm by (Fraigniaud et al., 2016; Barenboim et al., 2022; Maus and Tonoyan, 2022). Finally, we follow the key idea of the algorithm by Panconesi and Srinivasan (1992) to turn ${(\Delta+1)}$-coloring into a $\Delta$-coloring, and again we will need some advice to make this part efficient.
1.7. Contribution 6: vertex 3-coloring
@@
In the distributed setting, this does not work: $2$-coloring in bipartite graphs is still a global problem. Nevertheless, in Section 7 we show that it is still doable:
@@
1.8. Key technique: composability framework
@@
It turns out that for many problems, it is easier to work with advice schemas in which only a few nodes carry advice bits, but they may carry many bits of advice. In Definition 2 we give the formal definition of such a schema, and in Definition 4 we give the formal definition of composable schemas, which satisfy the additional property that the ratio between the total amount of bits held by the nodes, and the total amount of nodes, can be made arbitrarily small in every large-enough neighborhood.
While the definition is a bit technical, it has two key properties, which we discuss in more detail in Section 9:
  1. (1)
@@ 1.9. Open questions
  1. (1)
    Our negative result in Section 8 assumes the Exponential-Time Hypothesis. Is it possible to prove an unconditional lower bound without such assumptions? Can we exhibit a concrete LCL problem $\Pi$ that is unconditionally hard?
  2. (2)
@@
  3. (3)
    We also conjecture that for a sufficiently large $d$, vertex $d$-coloring in $d$-colorable graphs cannot be encoded with one bit per node. Is this true?
  4. (4)
@@
2. Related work
In (Feuilloley et al., 2021), the authors consider the framework of Proof Labeling Schemes (PLS). In this framework, the goal is to understand whether a given graph satisfies some predicate in a setting where we are allowed to assign labels (or certificates) to the nodes in order to help them achieve this goal. More precisely, in this framework there is a prover and a verifier: the prover is a centralized entity that assigns labels to the nodes, while the verifier is a distributed algorithm that in $T$ rounds is able to verify the validity of the collection of labels. If the predicate we want to investigate holds true, then there must exist an assignment of labels (or certificates) such that each node, in $T$ rounds, “accepts” the given labeling (i.e., the proof is valid in all $T$-hop neighborhoods); while if the predicate we want to investigate does not hold, then for any given labeling assignment there must exist a node that rejects it (i.e., that in its $T$-hop neighborhood notices that there is something wrong). The authors study the tradeoff between the size of the labels given at each node and the round-complexity of the verifier. For example, they show that, for all Proof Labeling Schemes on specific graph classes (such as trees, cycles, multidimensional grids, toruses), if some predicate has a Proof Labeling Scheme of $\beta$ bits using a $1$-round verifier, then it has a Proof Labeling Scheme of $O(\beta/T)$ bits using a $T$-round verifier (i.e., the size of the labels scales linearly with the runtime of the verifier algorithm). Moreover, they show that, on general graphs, if we consider all and only those predicates having a Proof Labeling Scheme such that $\beta$ bits are the same in all assigned labels, then the size of the uniform part of the labels can be reduced by a multiplicative factor of $\Omega(T)$. This gain is significant on graphs where the sizes of the neighborhoods increase a lot as a function of the distance. Finally, the authors focus on Proof Labeling Schemes for specific problems of interest, such as for proving that a given subset of edges is a spanning tree or a minimum-weight spanning tree, or for proving the diameter. They show lower and upper bounds on how much can the size of the labels scale with the runtime of the verifier. For example, for diameter, on the upper bound side, they show that for any $T$ in the range $[O(\log n),n]$ there exists a Proof Labeling Scheme with label-size of $O(n\log^{2}(n)/T)$ bits. For the same predicate, they show that for every $T$ in the range $[1,n/\log n]$, the size of the labels for proving the diameter of the graph is lower bounded by $\Omega(n/(T\log n))$ bits.
In (Ardévol Martínez et al., 2023), the authors consider the framework of Proof Labeling Schemes restricted to verifiers that run in exactly $1$ round, that is, each node has only access to its own label and the ones of its neighbors. The predicate that the authors study is $k$-colorability. While it is trivial to design a Proof Labeling Scheme for $2$-colorability by using a single bit per node and a $1$-round verifier, the authors show that this is not possible for $k$-colorability, where $k\geq 3$. For this problem, asymptotically tight bounds on the size of the labels are shown in (Bousquet et al., 2024). Note that the results shown in our paper imply that, for $k=3$, allowing $T(\Delta)$ rounds instead of $1$ round makes a huge difference, as in this case it is possible to verify $3$-colorability by using one bit per node.

### v2

@@
Abstract
Algorithms with advice have received ample attention in the distributed and online settings, and they have recently proven useful also in dynamic settings. In this work we study local computation with advice: the goal is to solve a graph problem $\Pi$ with a distributed algorithm in $T(\Delta)$ communication rounds, for some function $T$ that only depends on the maximum degree $\Delta$ of the graph, and the key question is how many bits of advice per node are needed.
Some of our results regard Locally Checkable Labeling problems (LCLs), which is an important family of problems that includes various coloring and orientation problems on finite-degree graphs. These are constraint-satisfaction graph problems that can be defined with a finite set of valid input/output-labeled neighborhoods.
Our main results are:
  1. 1.
    Any locally checkable labeling problem can be solved with only $1$ bit of advice per node in graphs with sub-exponential growth (the number of nodes within radius $r$ is sub-exponential in $r$; for example, grids are such graphs). Moreover, we can make the set of nodes that carry advice bits arbitrarily sparse. As a corollary, any locally checkable labeling problem admits a locally checkable proof with $1$ bit per node in graphs with sub-exponential growth.
  2. 2.
    The assumption of sub-exponential growth is complemented by a conditional lower bound: assuming the Exponential-Time Hypothesis, there are locally checkable labeling problems that cannot be solved in general with any constant number of bits per node.
  3. 3.
    In any graph we can find an almost-balanced orientation (indegrees and outdegrees differ by at most one) with $1$ bit of advice per node, and again we can make the advice arbitrarily sparse. As a corollary, we can also compress an arbitrary subset of edges so that a node of degree $d$ stores only $d/2+2$ bits, and we can decompress it locally, in $T(\Delta)$ rounds.
  4. 4.
@@ Abstract
  5. 5.
    In any $3$-colorable graph, we can find a $3$-coloring with $1$ bit of advice per node. As a corollary, in bounded-degree graphs there is a locally checkable proof that certifies $3$-colorability with $1$ bit of advice per node, while prior work shows that this is not possible with a proof labeling scheme (PLS), which is a more restricted setting where the verifier can only see up to distance $1$.
Our work shows that for many problems the key threshold is not whether we can achieve $1$ bit of advice per node, but whether we can make the advice arbitrarily sparse. To formalize this idea, we develop a general framework of composable schemas that enables us to build algorithms for local computation with advice in a modular fashion: once we have (1) a schema for solving $\Pi_{1}$ and (2) a schema for solving $\Pi_{2}$ assuming an oracle for $\Pi_{1}$, we can also compose them and obtain (3) a schema that solves $\Pi_{2}$ without the oracle. It turns out that many natural problems admit composable schemas, all of them can be solved with only $1$ bit of advice, and we can make the advice arbitrarily sparse.
keywords:
Distributed graph algorithms, LOCAL model, computation with advice, locally checkable labeling problems, proof labeling schemes, locally checkable proofs, graph coloring, exponential-time hypothesis
1 Introduction
@@
While computation with different forms of advice has been explored in a wide range of distributed settings [20, 23, 14, 32, 13, 36, 47, 49, 25, 37, 52, 21, 40, 48, 26], there is hardly any prior work on solving classic local graph problems. A rare example of prior work is [22] from 2007, which studied the question of how much advice is necessary to break Linial’s [45] lower bound for coloring cycles.
We initiate a systematic study of exactly how much advice is needed in the context of a wide range of graph problems. As we will see in this work, the exploration of the advice complexity of graph problems opens up connections with many other topics—it is linked with distributed proofs [41, 42, 43, 44, 33, 17], distributed decompression, the notion of order-invariant algorithms [51], and also with the exponential-time hypothesis [38] in computational complexity theory.
1.1 Local computation with advice
@@
We note that on graphs that do not admit a solution to $\Pi$, Algorithm $\mathcal{A}$ is allowed to behave arbitrarily. For instance, when we consider the $3$-coloring problem, Algorithm $\mathcal{A}$ may produce an arbitrary output on graphs that are not $3$-colorable.
@@
Our work shows that for many problems the key threshold is not whether we can achieve $1$ bit of advice per node, but whether we can make the advice arbitrarily sparse, that is, make the ratio between $1$s and 0s assigned to the nodes of the graph to be an arbitrarily small constant. This is particularly useful, as it enables us to compose multiple sparse advice schemes, so that it suffices to use just one bit per node in total (we will elaborate on this in Section˜1.7). Hence, a large part of this paper addresses the following question:
Which problems admit arbitrarily sparse advice?
@@
The notion of the sparsity of the advice is discussed in more detail later in the paper. In this context, our paper shows that some problems can be solved with arbitrarily sparse advice. On the other hand, we also show that assuming the Exponential-Time Hypothesis, for any constant $c$, there exist problems that cannot be solved with $c$ bits of advice. Finally, there are some problems like 3-coloring that can be solved with 1 bit of advice, but where it is not clear whether it can be solved with arbitrarily sparse advice. We discuss all of those points in more detail in the remaining part of this section.
1.2 Contribution 1: LCLs in bounded-growth graphs
@@
Locally checkable labeling problems (LCL), first introduced by Naor and Stockmeyer [51] in the 1990s, are one of the most extensively studied families of problems in the theory of distributed graph algorithms. These are graph problems that can be specified by giving a finite set of valid local neighborhoods. Many key problems such as vertex coloring, edge coloring, maximal independent set, maximal matching, sinkless orientation, and many other splitting and orientation problems are examples of LCLs, at least when restricted to bounded-degree graphs. Thanks to the extensive research effort since 2016, we now understand very well the landscape of all LCL problems and their computational complexities across different models of distributed computing [2, 4, 10, 28, 3, 19, 55, 9, 11, 27, 5].
@@
In Appendix˜A, we design a schema that allows us to solve any LCL problem with just one bit of advice in graphs with a sub-exponential growth (the number of nodes in a radius-$r$ neighborhood is sub-exponential in $r$):
@@
Any LCL problem can be solved with $1$ bit of advice per node in sub-exponential growth graphs.
Furthermore, we show that the encoding can be made arbitrarily sparse. Note that e.g. grids have polynomial growth while e.g. regular trees have exponential growth, so the result is applicable in grids but not in trees.
1.2.1 Application: locally checkable proofs in bounded-growth graphs
@@
One prominent application of this result is its connection with distributed proofs [41, 42, 43, 44, 17], and in particular with locally checkable proofs [33]. Consider any LCL $\Pi$. Assume that our task is to prepare a distributed proof that shows that in a graph $G$ there exists a feasible solution of $\Pi$ (for example, if $\Pi$ is the task of 10-coloring, then the task is to certify that the chromatic number of $G$ is at most 10). Now if $G$ has sub-exponential growth, we can use our result from Appendix˜A to prepare a $1$-bit advice that enables the algorithm to find a solution of $\Pi$. Our advice is the proof: to verify it, we simply try to recover a solution with the help of the advice, and then check that the output is feasible in all local neighborhoods (recall that $\Pi$ is locally checkable). We obtain the following corollary:
@@
Note that this is not a proof labeling scheme as defined in [41, 42, 43, 44], as the verifier running at node $u$ may need to see more than just the identifier and the proof label of $u$ and the proof labels of $u$’s immediate neighbors. However, in bounded-degree graphs it is a locally checkable proof (LCP) as defined in [33]; for a fixed $\Delta$ the verification radius is a constant $T(\Delta)$.
So to summarize, if we can solve some LCL problem $\Pi$ with $b$ bits of advice per node, then we also have an LCP for the graph property “G admits a feasible solution to $\Pi$” with $b$-bit proofs per node. The converse is not true: Consider the LCL problem $\Pi$ that encodes the task “orient edges so that each node has indegree equal to outdegree”. Now to prove “$G$ admits a feasible solution to $\Pi$” one can use a 0-bit LCP, where even-degree nodes accept and odd-degree nodes reject. However, this LCP does not help us at all if we would like to solve $\Pi$ with the help of advice. In this sense distributed computation with advice is a harder problem than local proofs.
It is also good to note that one can have LCPs for graph properties that are not of the form “G admits a feasible solution to some LCL $\Pi$.” For example, planarity is such a property. Such LCPs are (to our knowledge) not directly connected with computation with local advice.
1.3 Contribution 2: LCLs in general graphs
At this point a natural question is whether the assumption about bounded growth is necessary. Could we solve all LCL problems in all graphs with $1$ bit of advice? In Section˜4 we show that the answer is likely to be no:
@@ 1.3 Contribution 2: LCLs in general graphs
Hence, we need to show that $\mathcal{A}$ can be made cheap to simulate. The key ingredient is the following technical result, which we prove using a Ramsey-type argument that is inspired by the proof of Naor and Stockmeyer [51]:
@@
The key point here is that (for bounded-degree graphs) $\mathcal{A}^{\prime}$ can be represented as a finite lookup table; hence the simulation of $\mathcal{A}^{\prime}$ is cheap, and we can finally make a formal connection to the Exponential-Time Hypothesis. We refer to Section˜4 for more details.
1.4 Contribution 3: balanced orientations
In Appendix˜B we move on to a specific graph problem: we study the task of finding balanced and almost-balanced orientations. The goal is to orient the edges so that for each node indegree and outdegree differ by at most $1$. This is a hard problem to solve in a distributed setting, while slightly more relaxed versions of the problem admit efficient (but not constant-time) algorithms [30].
@@ 1.4 Contribution 3: balanced orientations
1.4.1 Application: distributed decompression
Equipped with the advice schema for solving almost-balanced orientations, we can now make a formal connection to what we call distributed decompression. Here the task is to encode some graph labeling so that it can be decompressed locally (in $T(\Delta)$ rounds).
@@ 1.4.1 Application: distributed decompression
1.5 Contribution 4: vertex $\Delta$-coloring
In Appendix˜C, we study the problem of $\Delta$-coloring graphs of maximum degree $\Delta$:
@@ 1.5 Contribution 4: vertex $\Delta$-coloring
Our schema for encoding $\Delta$-colorings consists of three steps. First, we compute a vertex coloring with $O(\Delta^{2})$ colors, with the help of advice. Then we reduce the number of colors down to $\Delta+1$, using the algorithm by [24, 6, 46]. Finally, we follow the key idea of the algorithm by Panconesi and Srinivasan [53] to turn $(\Delta+1)$}-coloring into a $\Delta$-coloring, and again we will need some advice to make this part efficient.
1.6 Contribution 5: vertex 3-coloring
@@
In the distributed setting, the trivial solution does not work: $2$-coloring in bipartite graphs is still a global problem. Nevertheless, in Appendix˜D we show that $3$-coloring is still doable with $1$ bit of advice:
@@
1.6.1 Application: locally checkable proofs for 3-coloring
Following the same idea as in Section˜1.2.1, we can now also certify that a graph is $3$-colorable with a proof that only takes $1$ bit per node. Furthermore, the verifier only needs to see up to distance $T(\Delta)$, and hence if $\Delta=O(1)$, this is a locally checkable proof in the sense of [33]:
There is a locally checkable proof that certifies $3$-colorability with $1$ bit per node in graphs of maximum degree $\Delta=O(1)$.
Now it is interesting to compare this with proof labeling schemes (PLS). In essence, a PLS is an LCP in which the verifier only sees the identifier of the present node, and the proof labels within radius $1$. The above result provides a $1$-bit LCP but not a $1$-bit PLS.
This connects directly with the work done by Ardévol Mart´ınez, Caoduro, Feuilloley, Narboni, Pournajafi, and Raymond [1] and Bousquet, Feuilloley, and Zeitoun [7], which study the same question: how many bits per node are needed to certify $k$-colorability. Here Bousquet et al. [7] shows that $\Theta(\log k)$ bits per node are necessary for a PLS that certifies $k$-colorability, and [1] shows that in particular $3$-colorability cannot be certified with $1$ bit per node with any PLS (while $2$ bits per node is trivial). Our work complements the latter result, and provides a separation between PLSs and LCPs in this setting: now we know that while $1$ bit per node does not suffice to certify $3$-colorability with any PLS, it is sufficient for an LCP (with the caveat that we need to be in a bounded-degree graph).
1.7 Key technique: composability framework
@@
It turns out that for many problems, it is easier to work with advice schemas in which only a few nodes carry advice bits, but they may carry many bits of advice. In Definition˜3.2 we give the formal definition of such a schema, and in Definition˜3.4 we give the formal definition of composable schemas, which satisfy the additional property that the ratio between the total number of bits held by the nodes, and the total number of nodes, can be made arbitrarily small in every large-enough neighborhood.
While the definition is a bit technical, it has two key properties, which we discuss in more detail in Appendix˜E:
  1. 1.
@@ 1.8 Open questions
  1. 1.
    Our negative result in Section˜4 assumes the Exponential-Time Hypothesis. Is it possible to prove an unconditional lower bound without such assumptions? Can we exhibit a concrete LCL problem $\Pi$ that is unconditionally hard?
  2. 2.
@@
  3. 3.
    We also conjecture that for a sufficiently large $d$, vertex $d$-coloring in $d$-colorable graphs cannot be encoded with one bit per node. Is this true? This is closely linked with the discussion on the trade-off conjecture in e.g. [1, 7].
  4. 4.
@@
2 Additional related work
Computing with advice is not really a well-defined area of research, as one can consider many kinds of advice and various models of computations. In this paper, we focus on existential advice and the distributed LOCAL model: the advice is provided by an all-knowing oracle, and it is designed to enable algorithms with low locality (i.e., low time complexity). More broadly, one can consider a variant in which the restriction on the size of advice is more relaxed, but it needs to be computable by an oracle with some limitations (realizable in some model of computation), which then can be used as a building block of algorithms.
Advice and locality in distributed computing.
Feuilloley, Fraigniaud, Hirvonen, Paz, and Perry [18] consider the framework of Proof Labeling Schemes (PLS). In this framework, there is a prover and a verifier: the prover is a centralized entity that assigns labels to the nodes, while the verifier is a distributed algorithm that in $T$ rounds is able to verify the validity of the collection of labels. If the predicate we want to investigate holds true, then there must exist an assignment of labels (or certificates) such that each node, in $T$ rounds, “accepts” the given labeling; while if the predicate we want to investigate does not hold, then for any given labeling assignment there must exist a node that rejects it. The authors study the tradeoff between the size of the labels given at each node and the round-complexity of the verifier. In their work, they, too, discover how this tradeoff is remarkably different depending on the growth rate of the graph.